OpenStudy (anonymous):

Evaluate x^-2 for x = -3

OpenStudy (anonymous):

\[x ^{-2} = 1/x ^{2}\] \[= 1/ (-3)^{2} =???\]

OpenStudy (anonymous):

dude i have no idea what you just did lol

OpenStudy (anonymous):

when you see a negative sign on the exponent, then it means you get the reciprocal, in this case the exponent is -2, so you get the reciprocal of the square of that number, got it?

OpenStudy (anonymous):

yeah i just know the anser should be a fraction lol

OpenStudy (anonymous):

so what's your answer?

OpenStudy (anonymous):

i dont know thats why im on here...

OpenStudy (anonymous):

but i've already given the answer: \[1/(-3)^{2} = 1/9\]